# Stealthy Patch-Wise Backdoor Attack in 3D Point Cloud via Curvature Awareness

### v1

@@ A.2 Effect of the Poisoning Rate $\rho$
A.3 Effect of the Selected Patch Number $m$
In this section, we experiment with injecting the spectral trigger into varying numbers $m$ of selected patches during the training process. The experiments are conducted on ModelNet40 using PointNet++. As shown in Table 6, as $m$ increases from 1 to 16, the Chamfer Distance (CD) gradually decreases, indicating improved stealthiness. This suggests that selecting more patches within a reasonable range helps distribute perturbations more naturally, enhancing imperceptibility. However, when $m$ further increases to 32, where almost all points are perturbed, CD rises to 0.12, indicating that excessive modifications start to introduce noticeable distortions, reducing stealthiness. Based on these observations, we set $m=16$ as the default configuration, as it achieves the optimal balance between imperceptibility (lowest CD), strong attack effectiveness, and stable benign accuracy.
@@ A.3 Effect of the Selected Patch Number $m$
Appendix B Visualization on More Poisoned Samples

### v2

@@ 4.5 Resistance to Defense Methods
4.6 Ablation Study
Effect of Patch Selection Strategies. In our default setting, we select the 16 patches with the highest patch imperceptibility scores (PIS), denoted as H-PSI. To evaluate the effectiveness of this selection strategy, we compare H-PSI with three alternative patch selection strategies: (1) L-PSI: Select the patches with the lowest PIS; (2) Random: Randomly select the patches; (3) FPSP: Use the FPS algorithm to sample the same number of points as H-PSI. We evaluate their average performance on ModelNet40. As shown in Tab. 7, the H-PSI strategy achieves the highest attack effectiveness and the strongest stealthiness. This result underscores the advantage of selecting high-imperceptibility patches for local spectral trigger injection.
Effect of Selected Patch Number $m$. In this section, we experiment with injecting the spectral trigger into varying numbers $m$ of selected patches during the training process. The experiments are conducted on ModelNet40 using PointNet++. As shown in Tab. 8, as $m$ increases from 1 to 16, the CD value gradually decreases, indicating improved stealthiness. This suggests that selecting more patches within a reasonable range helps distribute perturbations more naturally, enhancing imperceptibility. However, when $m$ further increases to 32, where almost all points are perturbed, CD rises to 0.12, indicating that excessive modifications start to introduce noticeable distortions, reducing stealthiness. Based on these observations, we set $m=16$ as the default configuration, as it achieves the optimal balance between imperceptibility (lowest CD), strong attack effectiveness, and benign accuracy.
5 Conclusion
In this paper, we propose SPBA, a novel backdoor attack method against 3D point cloud models. Unlike existing methods that rely on sample-wise trigger injection, which introduces noticeable geometric distortions, SPBA adopts a local patch-wise spectral trigger that selectively perturbs geometrically complex regions, demonstrating superior stealthiness, efficiency and effectiveness. Additional experiments validate the strong resistance of SPBA to existing defense methods. Experimental results highlight the effectiveness of patch-wise triggers as a promising direction for stealthy and adaptive backdoor attacks in 3D point clouds.
Appendix
Appendix A Deformation by Spectral Perturbation
As illustrated in Figure 4, low-spectrum perturbations cause significant deformation to the patch of empennage, while high-spectrum perturbations introduce excessive outliers. Perturbations applied within intermediate spectral bands mitigate outliers while preserving the overall shape, presenting a key challenge for the optimal design of spectral triggers. To address this, we propose an optimizable spectral trigger $\xi$ that strikes a balance between attack effectiveness and stealthiness.
Figure 4: Comparison of original and perturbed patches. The second row shows the spectral features of the original patch and spectral perturbations across different spectrum bands.
Appendix B Resistance to the Gradient-based Method
As many studies on backdoor attacks assess trigger stealthiness using gradient-based methods [22, 17], we employ a point cloud saliency map [35] to identify salient points that contribute most to the model’s decision-making. Specifically, 40 points with the highest saliency scores are marked in red. As illustrated in Figure 5, the spherical trigger of PointBA-I is easily detected. In contrast, our proposed SPBA exhibits a saliency distribution similar to the clean model, indicating enhanced stealthiness and reduced detectability.
Appendix C Ablation Study
C.1 Effect of the Target Class
@@
Figure 7: Effect of the target class selection.
C.2 Effect of the Poisoning Rate
Figure 8: The effect of the poisoning rate.
Appendix D Visualization on More Poisoned Samples